OpenStudy (anonymous):

How do I solve this quadratic equation by factoring: 9x^2 - 25 =0

OpenStudy (anonymous):

(3x+5)(3x-5) (a+b)(a-b)=a^2 - b^2

OpenStudy (anonymous):

how?

OpenStudy (anonymous):

(3x)^2 -(5)^2=(3x+5)(3x-5)

jhonyy9 (jhonyy9):

\[(3x-5)(3x+5)=9x ^{2}-25\]